OpenStudy (anonymous):

Solving equations with Radicals:- sqr x^2+15x-6=x

OpenStudy (amistre64):

you need to use (...)s to indicate what is spose to be under the radical sign. sqrt(x^2 +15x -6) = x would be appropriate, otherwise we got no idea what you need from it

OpenStudy (anonymous):

I'm assuming you mean: \[\sqrt{x^2 + 15x - 6}=x\] In that case: square both sides: \[x^2+15x-6 = x^2\] \[15x-6=0\] \[15x=6\] \[x=6/15\]

OpenStudy (dumbcow):

usually the idea is to get rid of the radical by squaring both sides

OpenStudy (anonymous):

oh yes u r right

OpenStudy (dumbcow):

with radicals, always check for extraneous solutions :) plug it back in and make sure it works

OpenStudy (amistre64):

sqrt(x^2) +15x -6 = x would be: x +15x -6 -x = 0 15x -6 = 0 15x = 6 x = 6/15 = 2/5